# Integrating Dynamical Systems Modeling with Spatiotemporal scRNA-seq Data Analysis

### v1

@@ 5.1 Bridging Discrete and Continuous Dynamics Modeling
One interesting topic to explore is building connections between discrete dynamic models (e.g. Markov Chain) with continuous differential equations when dealing with scRNA-seq data. In CellRank (Lange et al., 2022; Weiler et al., 2024), the output from continuous models could help to refine the random walk on the data point cloud by introducing various kernels. Let the spliced RNA counts of cells $i$ and $j$ be $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$, respectively. The Gaussian Kernel is defined as
$d_{g}(\mathbf{s}_{i},\mathbf{s}_{j})=\exp\left(-\frac{\|\mathbf{s}_{i}-\mathbf{s}_{j}\|^{2}}{\sigma^{2}}\right),$
Let the position vector from cell $i$ to cell $j$ be $\mathbf{\delta}_{ij}=\mathbf{s}_{j}-\mathbf{s}_{i}$ and $v_{i}$ denotes the estimated RNA velocity of cell $i$. Then, the three velocity kernels can be introduced as
  • Cosine Kernel: $v_{\cos}(\mathbf{s}_{i},\mathbf{s}_{j})=g\left(\cos(\mathbf{\delta}_{ij},\mathbf{v}_{i})\right)$,
  • Correlation Kernel: $v_{\text{corr}}(\mathbf{s}_{i},\mathbf{s}_{j})=g\left(\mathrm{corr}(\mathbf{\delta}_{ij},\mathbf{v}_{i})\right)$,
  • Inner Product Kernel: $v_{\text{ip}}(\mathbf{s}_{i},\mathbf{s}_{j})=g\left(\mathbf{\delta}_{ij}^{T}\mathbf{v}_{i}\right)$.
Here $g$ is a bounded, positive, monotonic increasing function such as an exponential function. In CellRank, the actual transition kernel can combine these two parts either by weighted summation or multiplication. For example, the kernel used in the original RNA Velocity method is:
$\mathrm{Ker}(\mathbf{s}_{i},\mathbf{s}_{j})=\lambda v_{\cos}(\mathbf{s}_{i},\mathbf{s}_{j})+(1-\lambda)d_{g}(\mathbf{s}_{i},\mathbf{s}_{j}),$
where $\lambda$ is a weighting coefficient. The Markov chain transition matrix is constructed as:
$p_{ij}=\frac{\mathrm{Ker}(\mathbf{s}_{i},\mathbf{s}_{j})}{\sum_{j}\mathrm{Ker}(\mathbf{s}_{i},\mathbf{s}_{j})}.$
CellRank2 (Weiler et al., 2024) provides more flexible options for the transition kernel and incorporates prior knowledge. For example, if pseudotime is known in advance, it can be used to adjust the transition kernel:
$\mathrm{Ker}_{\text{adj}}(\mathbf{s}_{i},\mathbf{s}_{j})=\mathrm{Ker}(\mathbf{s}_{i},\mathbf{s}_{j})f(\Delta t_{ij}),$
@@
Another direction is to analyze the theoretical convergence of discrete dynamics as the number of data points tends to infinity. A well-known example is the study of the continuum limit of diffusion map random walk (Coifman et al., 2008), stating that the random walk induced by the Gaussian kernel would converge to the dynamics of the Fokker-Planck equation. When considering growth, the directed random walk defined by PBA would converge to (2.3). Interestingly, (Zhou et al., 2021b) also proves that the coarse-grained transition probabilities yield the continuum limit of transition rate among attractors (3.6), therefore validating the rationale of MuTrans.
@@
Once such a theoretical connection is built, new theoretical insights could be drawn toward the algorithm design. For instance, (Li et al., 2020) systematically investigates the continuous limit of RNA-velocity-induced random walk kernels. For example, if the transition kernel is $\mathrm{Ker}(\mathbf{s}_{i},\mathbf{s}_{j})=d_{g}(\mathbf{s}_{i},\mathbf{s}_{j})\cdot v_{\cos}(\mathbf{s}_{i},\mathbf{s}_{j}),$ the corresponding ODE yields the desired streamlined equation that correctly reveals the vector field directionality $\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}=\frac{\mathbf{v}}{\|\mathbf{v}\|}.$ Meanwhile, for $\mathrm{Ker}(\mathbf{s}_{i},\mathbf{s}_{j})=d_{g}(\mathbf{s}_{i},\mathbf{s}_{j})\cdot v_{\text{corr}}(\mathbf{s}_{i},\mathbf{s}_{j})$, the corresponding ODE is $\frac{\mathrm{d}\mathbf{x}}{\mathrm{d}t}=\frac{\mathcal{P}_{\mathbf{1}}\mathbf{v}}{\|\mathcal{P}_{\mathbf{1}}\mathbf{v}\|},$ where $\mathcal{P}$ is the projection operator defined as $\mathcal{P}_{n}\mathbf{x}=(I-\mathbf{\hat{n}}\otimes\mathbf{\hat{n}})\cdot\mathbf{x},\quad\mathbf{\hat{n}}=\frac{\mathbf{n}}{\|\mathbf{n}\|},\quad\mathbf{1}=(1,1,\cdots,1)^{T}.$ which indicates that the correlation kernel might alter both the direction and magnitude of RNA velocity in the continuous limit.
@@
Some approaches also use discrete graphs to represent the geometric structure of data (Hetzel et al., 2021). Graphdynamo (Zhang et al., 2023) and Graphvelo (Chen et al., 2024) propose a method that leverages geometric structure to correct RNA Velocity. They assume that the cell data points $\mathbf{x}_{i}$ lie on a low-dimensional manifold embedded in a high-dimensional space (in classical mechanics, this is known as the “configuration manifold”) and that each cell’s RNA Velocity vector lies in the tangent space $T_{x}\mathcal{M}$ at the point $\mathbf{x}_{i}$ on the manifold. Let $\mathbf{\delta}_{ij}$ denote the displacement vector from cell $i$ to its neighboring cell $j$. With a sufficient number of such $\mathbf{\delta}_{ij}$, one can construct a non-orthogonal normalized basis for $T_{x}\mathcal{M}$. Thus, the RNA Velocity vector in $T_{x}\mathcal{M}$ can be expressed as $\mathbf{v}_{\parallel}(\mathbf{x}_{i})=\sum_{j\in\mathcal{N}_{i}}\phi_{ij}\mathbf{\delta}_{ij}$ The coefficients of the linear combination, $\mathbf{\phi}_{i}=\{\phi_{ij}\mid j\in\mathcal{N}_{i}\}$, are determined by minimizing the following loss
$\mathcal{L}(\phi_{i})=\|\mathbf{v}_{i}-\mathbf{v}_{\parallel}(\mathbf{x}_{i})\|^{2}-b\cos(\mathbf{\phi}_{i},\mathbf{\phi}_{i}^{\text{corr}})+\lambda\|\mathbf{\phi}_{i}\|^{2}$
@@
Furthermore, the population dynamics in the feature space can be transferred to the dynamics on the graph. The unbalanced Fokker-Planck equation (2.3) could be generalized to a graph, such that the mass evolution at node $i$ is given by:
@@
5.2 Cell-cell Interaction
For temporally resolved single-cell RNA-seq data, previous modeling approaches have typically developed in the continuous space of $\mathbb{R}^{d}$. However, the data inherently exists within a discrete space. Therefore, an interesting question is how to construct continuous cellular dynamics from a discrete space, e.g., a graph (Gandrillon and Stumpf, 2021). In (Jiang et al., 2022), it proposes GraphFP, a graph Fokker-Planck equation-based method to model cellular dynamics. Consider the data are collected at $T$ time points $\{t_{1},t_{2}\cdots t_{T}\}$, at each point $t_{i}$ there are $n_{i}$ cells are measured with the corresponding expression vector $\mathbf{x}_{1}^{(t_{i})},\mathbf{x_{2}}^{(t_{i})},\cdots,\mathbf{x}_{n_{i}}^{t_{i}}\in\mathbb{R}^{d}$. Thus there are $N=\sum_{i=1}^{T}n_{i}$ cells in total. We assume the data can be clustered or annotated into $M$ cell types. Next, they consider constructing a cell state transition graph ${G}=(V,E)$, where each vertex in $V$ represents a cell type and each edge $\{i,j\}$ in $E$ means the cell type $i$ can transit to cell type $j$.
Unlike the previous methods considering probability distribution in $\mathbb{R}^{d}$, now they consider the probability distribution on graph ${G}$. Suppose there are $M$ vertices in graph ${G}$, we consider the probability simplex supported on all vertices of ${G}$
$\mathcal{P}({G})=\{\mathbf{p}(t)=(p_{i}(t))_{i=1}^{M}\mid\sum_{i=1}^{M}p_{i}(t)=1,p_{i}(t)\geq 0\}.$
@@ 5.2 Cell-cell Interaction
The aim is also to transport the distribution from $\mathbf{p}_{0}$ to $\mathbf{p}_{1}$ on $G$, satisfying certain least action principles. So like the continuous space, we need to define the Fokker-Planck equation and the Wasserstein distance on graph $G$. First, we can define a free energy $\mathcal{F}:\mathcal{P}(G)\rightarrow\mathbb{R}$, then the Fokker-Planck equation can be defined as follows:
$\frac{\mathrm{d}p_{i}(t)}{\mathrm{d}t}=\sum_{j\in\mathcal{N}(i)}\left(\frac{\partial\mathcal{F}(\mathbf{p})}{\partial p_{j}}-\frac{\partial\mathcal{F}(\mathbf{p})}{\partial p_{i}}\right)g_{ij}(\mathbf{p}),$
@@
where $\mathcal{N}(i)$ is adjacency set of vertex $i$ and $g_{ij}(\mathbf{p})$ satisfy certain definitions (Jiang et al., 2022). Next, the discrete L2-Wasserstein distance on graph $G$ between $\mathbf{p}_{0},\mathbf{p}_{1}\in\mathcal{P}(G)$ can be defined as
$\mathcal{W}_{2,G}^{2}(\mathbf{p}_{0},\mathbf{p}_{1})=\inf_{\mathcal{F}}\frac{1}{2}\int_{0}^{1}\sum_{i,j\in E}\left(\frac{\partial\mathcal{F}(\mathbf{p})}{\partial p_{j}}-\frac{\partial\mathcal{F}(\mathbf{p})}{\partial p_{i}}\right)^{2}g_{ij}(\mathbf{p}).$
@@
Next, the target is to find the minimum energy path. In (Jiang et al., 2022), they parameterize $\mathcal{F}$ by a linear energy form
$\mathcal{F}(\mathbf{p}\mid\mathbf{\Phi},\mathbf{W})$ $=\mathcal{V}(\mathbf{p})+\mathcal{W}(\mathbf{p})+\beta\mathcal{H}(\mathbf{p}),$
@@
$=\mathbf{p}^{T}\mathbf{\Phi}+\frac{1}{2}\mathbf{p}^{T}\mathbf{W}\mathbf{p}+\beta\sum_{i=1}^{n}p_{i}\log p_{i},$
where $\mathbf{\Phi}=(\Phi_{i})_{i=1}^{M}$, $\mathbf{W}=(w_{i,j})_{1\leq i,j\leq M}$ and $\beta\geq 0$ is a hyper-parameter. So after this parametrization, we hereinafter denote the parameters of the free energy as $\mathbf{\theta}=\{\mathbf{\Phi},\mathbf{W}\}$, and the goal is to find the parameter $\mathbf{\theta}$ such that
$\mathbf{\theta}^{*}=\arg\min_{\mathbf{\theta}}\quad\int_{t_{1}}^{t_{f}}\frac{1}{2}\sum_{\{i,j\}\in E}\left(\frac{\partial\mathcal{F}(\mathbf{p})}{\partial p_{i}}-\frac{\partial\mathcal{F}(\mathbf{p})}{\partial p_{j}}\right)^{2}\cdot g_{ij}(\mathbf{p}(t))dt,$
@@
$\frac{d\mathbf{p}(t)}{dt}$ $=\left(\sum_{j\in\mathcal{N}(i)}\left(\frac{\partial\mathcal{F}(\mathbf{p})}{\partial p_{j}}-\frac{\partial\mathcal{F}(\mathbf{p})}{\partial p_{i}}\right)g_{ij}(\mathbf{p}(t))\right)_{i=1}^{M}$
$\mathbf{p}(1)$ $=\mathbf{p}_{1}$
This problem can be solved by the adjoint method. Once these dynamics are solved, then we can use them to do some downstream analysis, e.g., cell-cell interaction, probability flow of cell types, and the potential energy (Jiang et al., 2022).
Looking ahead, important directions for future research include the expansion of the current framework to achieve single-cell resolution instead of cellular types, incorporating the matching of unnormalized distribution results from cell proliferation and death, and extending the model to spatial transcriptomics.

### v2

@@ 5.1 Bridging Discrete and Continuous Dynamics Modeling
One interesting topic to explore is building connections between discrete dynamic models (e.g., Markov chain) with continuous differential equations when dealing with scRNA-seq data. In CellRank Lange et al. (2022); Weiler et al. (2024), the output from continuous models could help to refine the random walk on the data point cloud by introducing various kernels. Let the spliced RNA counts of cells $i$ and $j$ be $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$, respectively. The Gaussian Kernel is defined as
$d_{g}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})=\exp\left(-\frac{\|\boldsymbol{s}_{i}-\boldsymbol{s}_{j}\|^{2}}{\sigma^{2}}\right),$
Let the position vector from cell $i$ to cell $j$ be $\boldsymbol{\delta}_{ij}=\boldsymbol{s}_{j}-\boldsymbol{s}_{i}$ and $v_{i}$ denotes the estimated RNA velocity of cell $i$. Then, the three velocity kernels can be introduced as
  • Cosine Kernel: $v_{\cos}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})=g\left(\cos(\boldsymbol{\delta}_{ij},\boldsymbol{v}_{i})\right)$,
  • Correlation Kernel: $v_{\text{corr}}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})=g\left(\mathrm{corr}(\boldsymbol{\delta}_{ij},\boldsymbol{v}_{i})\right)$,
  • Inner Product Kernel: $v_{\text{ip}}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})=g\left(\boldsymbol{\delta}_{ij}^{T}\boldsymbol{v}_{i}\right)$.
Here, $g$ is a bounded, positive, monotonic increasing function such as an exponential function. In CellRank Lange et al. (2022), the actual transition kernel can combine these two parts either by weighted summation or multiplication. For example, the kernel used in the original RNA velocity method is
$\mathrm{Ker}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})=\lambda v_{\cos}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})+(1-\lambda)d_{g}(\boldsymbol{s}_{i},\boldsymbol{s}_{j}),$
where $\lambda$ is a weighting coefficient. The Markov chain transition matrix is constructed as
$p_{ij}=\frac{\mathrm{Ker}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})}{\sum_{j}\mathrm{Ker}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})}.$
CellRank2 Weiler et al. (2024) provides more flexible options for the transition kernel and incorporates prior knowledge. For example, if pseudotime is known in advance, it can be used to adjust the transition kernel
$\mathrm{Ker}_{\text{adj}}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})=\mathrm{Ker}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})f(\Delta t_{ij}),$
@@
Another direction is to analyze the theoretical convergence of discrete dynamics as the number of data points tends to infinity. A well-known example is the study of the continuum limit of diffusion map random walk Coifman et al. (2008), stating that the random walk induced by the Gaussian kernel would converge to the dynamics of the Fokker–Planck equation. When considering growth, the directed random walk defined by PBA would converge to (3). Interestingly, ref. Zhou et al. (2021) also proves that the coarse-grained transition probabilities yield the continuum limit of transition rate among attractors (6), therefore validating the rationale of MuTrans.
@@
Once such a theoretical connection is built, new theoretical insights could be drawn toward the algorithm design. For instance, ref. Li et al. (2020) systematically investigates the continuous limit of RNA-velocity-induced random walk kernels. For example, if the transition kernel is $\mathrm{Ker}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})=d_{g}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})\cdot v_{\cos}(\boldsymbol{s}_{i},\boldsymbol{s}_{j}),$ the corresponding ODE yields the desired streamlined equation that correctly reveals the vector field directionality $\frac{\mathrm{d}\boldsymbol{x}}{\mathrm{d}t}=\frac{\boldsymbol{v}}{\|\boldsymbol{v}\|}.$ Meanwhile, for $\mathrm{Ker}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})=d_{g}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})\cdot v_{\text{corr}}(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$, the corresponding ODE is $\frac{\mathrm{d}\boldsymbol{x}}{\mathrm{d}t}=\frac{\mathcal{P}_{\boldsymbol{1}}\boldsymbol{v}}{\|\mathcal{P}_{\boldsymbol{1}}\boldsymbol{v}\|},$ where $\mathcal{P}$ is the projection operator defined as $\mathcal{P}_{n}\boldsymbol{x}=(I-\boldsymbol{\hat{n}}\otimes\boldsymbol{\hat{n}})\cdot\boldsymbol{x},\quad\boldsymbol{\hat{n}}=\frac{\boldsymbol{n}}{\|\boldsymbol{n}\|},\quad\boldsymbol{1}=(1,1,\cdots,1)^{T}.$ which indicates that the correlation kernel might alter both the direction and magnitude of RNA velocity in the continuous limit.
@@
Some approaches also use discrete graphs to represent the geometric structure of data Hetzel et al. (2021). Graphdynamo Zhang et al. (2023) and Graphvelo Chen et al. (2024) propose a method that leverages geometric structure to correct RNA velocity. They assume that the cell data points $\boldsymbol{x}_{i}$ lie on a low-dimensional manifold embedded in a high-dimensional space (in classical mechanics, this is known as the “configuration manifold”) and that each cell’s RNA velocity vector lies in the tangent space $T_{x}\mathcal{M}$ at the point $\boldsymbol{x}_{i}$ on the manifold. Let $\boldsymbol{\delta}_{ij}$ denote the displacement vector from cell $i$ to its neighboring cell $j$. With a sufficient number of such $\boldsymbol{\delta}_{ij}$, one can construct a non-orthogonal normalized basis for $T_{x}\mathcal{M}$. Thus, the RNA velocity vector in $T_{x}\mathcal{M}$ can be expressed as $\boldsymbol{v}_{\parallel}(\boldsymbol{x}_{i})=\sum_{j\in\mathcal{N}_{i}}\phi_{ij}\boldsymbol{\delta}_{ij}$ The coefficients of the linear combination, $\boldsymbol{\phi}_{i}=\{\phi_{ij}\mid j\in\mathcal{N}_{i}\}$, are determined by minimizing the following loss:
$\mathcal{L}(\phi_{i})=\|\boldsymbol{v}_{i}-\boldsymbol{v}_{\parallel}(\boldsymbol{x}_{i})\|^{2}-b\cos(\boldsymbol{\phi}_{i},\boldsymbol{\phi}_{i}^{\text{corr}})+\lambda\|\boldsymbol{\phi}_{i}\|^{2}$
@@
Furthermore, the population dynamics in the feature space can be transferred to the dynamics on the graph. The unbalanced Fokker–Planck Equation (3) could be generalized to a graph, such that the mass evolution at node $i$ is given by the following equation:
@@
5.2 Modeling Cell–Cell Interaction Dynamically
For temporally resolved single-cell RNA-seq data, previous modeling approaches have typically been developed in the continuous space of $\mathbb{R}^{d}$. However, the data inherently exist within a discrete space. In addition, incorporating cell–cell communication or interaction into these dynamics are important for constructing accurate spatiotemporal developmental landscapes and for advancing our understanding of complex biological systems Almet et al. (2021); Jin et al. (2021, 2025); Cang et al. (2023); Almet et al. (2024); Wada et al. (2021); Topolewski and Komorowski (2021). Therefore, an interesting question is how to construct continuous cellular dynamics from a discrete space of interacting cells, e.g., those represented by a graph Gandrillon and Stumpf (2021).
In  Jiang et al. (2022), it proposes GraphFP, a graph Fokker–Planck equation-based method to model cellular dynamics by explicitly considering cell interactions. Assume that data can be clustered or annotated into $M$ cell types, GraphFP constructs a cell state transition graph ${G}=(V,E)$, where each vertex in $V$ represents a cell type and each edge $\{i,j\}$ in $E$ means the cell type $i$ can transit to cell type $j$. Unlike other methods considering probability distribution in $\mathbb{R}^{d}$, GraphFP consider the probability distribution on graph ${G}$. Suppose there are $M$ vertices in graph ${G}$, consider the probability simplex supported on all vertices of ${G}$
$\mathcal{P}({G})=\{\boldsymbol{p}(t)=(p_{i}(t))_{i=1}^{M}\mid\sum_{i=1}^{M}p_{i}(t)=1,p_{i}(t)\geq 0\}.$
@@ 5.2 Modeling Cell–Cell Interaction Dynamically
The aim is also to transport the distribution from $\boldsymbol{p}_{0}$ to $\boldsymbol{p}_{1}$ on $G$, satisfying least action principles. Similar to the continuous space case, one need to define the Fokker–Planck equation and the Wasserstein distance on graph $G$. First, one can define a free energy $\mathcal{F}:\mathcal{P}(G)\rightarrow\mathbb{R}$, then the Fokker–Planck equation can be defined as follows:
$\frac{\mathrm{d}p_{i}(t)}{\mathrm{d}t}=\sum_{j\in\mathcal{N}(i)}\left(\frac{\partial\mathcal{F}(\boldsymbol{p})}{\partial p_{j}}-\frac{\partial\mathcal{F}(\boldsymbol{p})}{\partial p_{i}}\right)g_{ij}(\boldsymbol{p}),$
@@
where $\mathcal{N}(i)$ is the neighbor set of vertex $i$ and $g_{ij}(\boldsymbol{p})$ satisfy certain constraints that could constructed from $\boldsymbol{p}$ Jiang et al. (2022). Next, the discrete L2-Wasserstein distance on graph $G$ between $\boldsymbol{p}_{0},\boldsymbol{p}_{1}\in\mathcal{P}(G)$ can be defined as
$\mathcal{W}_{2,G}^{2}(\boldsymbol{p}_{0},\boldsymbol{p}_{1})=\inf_{\mathcal{F}}\frac{1}{2}\int_{0}^{1}\sum_{i,j\in E}\left(\frac{\partial\mathcal{F}(\boldsymbol{p})}{\partial p_{j}}-\frac{\partial\mathcal{F}(\boldsymbol{p})}{\partial p_{i}}\right)^{2}g_{ij}(\boldsymbol{p}).$
@@
Next, the target is to find the minimum energy path. In  Jiang et al. (2022), they parameterize $\mathcal{F}$ by a linear energy form
$\mathcal{F}(\boldsymbol{p}\mid\boldsymbol{\Phi},\boldsymbol{W})$ $=\mathcal{V}(\boldsymbol{p})+\mathcal{W}(\boldsymbol{p})+\beta\mathcal{H}(\boldsymbol{p}),$
@@
$=\boldsymbol{p}^{T}\boldsymbol{\Phi}+\frac{1}{2}\boldsymbol{p}^{T}\boldsymbol{W}\boldsymbol{p}+\beta\sum_{i=1}^{n}p_{i}\log p_{i},$
where $\boldsymbol{\Phi}=(\Phi_{i})_{i=1}^{M}$, $\boldsymbol{W}=(w_{i,j})_{1\leq i,j\leq M}$ represents the interactions among cell types, and $\beta\geq 0$ is a hyper-parameter. After this parametrization, one denotes the parameters of the free energy as $\boldsymbol{\theta}=\{\boldsymbol{\Phi},\boldsymbol{W}\}$, and the goal is to find the parameter $\boldsymbol{\theta}$ such that
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\quad\int_{t_{1}}^{t_{f}}\frac{1}{2}\sum_{\{i,j\}\in E}\left(\frac{\partial\mathcal{F}(\boldsymbol{p})}{\partial p_{i}}-\frac{\partial\mathcal{F}(\boldsymbol{p})}{\partial p_{j}}\right)^{2}\cdot g_{ij}(\boldsymbol{p}(t))dt,$
@@
$\frac{d\boldsymbol{p}(t)}{dt}$ $=\left(\sum_{j\in\mathcal{N}(i)}\left(\frac{\partial\mathcal{F}(\boldsymbol{p})}{\partial p_{j}}-\frac{\partial\mathcal{F}(\boldsymbol{p})}{\partial p_{i}}\right)g_{ij}(\boldsymbol{p}(t))\right)_{i=1}^{M}$
$\boldsymbol{p}(1)$ $=\boldsymbol{p}_{1}$
This problem can be solved by the adjoint method. Once these dynamics are solved, one can then use them for downstream analysis, e.g., cell–cell interaction, probability flow of cell types, and the potential energy Jiang et al. (2022).
Looking ahead, we anticipate that important future directions based on GraphFP include the expansion of the current framework to achieve single-cell resolution instead of cellular types, incorporating the matching of unnormalized distribution results from cell proliferation and death, and extending the model to spatial transcriptomics.